OpenStudy (luigi0210):

Previous Question:

OpenStudy (luigi0210):

\(f'(x)=2x^3\), find f. Given the tangent 2x+y=0 First the general anti-derivative: \(F(x)=\frac{1}{2}x^4+c\) Got that \(y=-2x\), slope \(m=-2\), or f'(x)=-2\) right? So \(-2=2x^3\) and solving for x.. \(-1=x\). Plug that into the tangent \(2(-1)+y=0\) and get \(y=2\), giving the points \((-1, 2)\).

OpenStudy (luigi0210):

Finishing by plugging it back into \(F(x)=\frac{1}{2}x^4\) so.. \(2=\frac{1}{2}(-1)^4+C\) So \(C=1.5\)?

hartnn (hartnn):

seems correct to me :)

OpenStudy (luigi0210):

Alright, thank you hartnn :D